OpenStudy (anonymous):

(3/4)-(5/6) sos me pls step by step

OpenStudy (lgbasallote):

\[\frac 34 - \frac 56 \implies \frac{18}{24} - \frac{20}{24}\] does that help?

OpenStudy (anonymous):

how did u get 18/24 - 20/24?

OpenStudy (lgbasallote):

i took the LCD of 6 and 4...

OpenStudy (lgbasallote):

i just multiplied them...6 x 4 is 24 so i made the denominators into 24

OpenStudy (lgbasallote):

then to make 3/4 into a fraction 24/4 = 6 6x 3 = 18 hence 18/24

OpenStudy (anonymous):

and then i simplify?

OpenStudy (lgbasallote):

first subtract 18/24 - 20/24 then you simplify

OpenStudy (anonymous):

awesome